# arXiv:2602.22626v2[cond-mat.str-el] 02 Mar 2026

### v2

@@ Abstract
From this perspective, Hamiltonian energy eigenmodes describe conservative reversible dynamics, whereas relaxation spectra encode the irreversible organization governing collective behavior near criticality. Strongly correlated quantum matter near dynamical critical points may thus be more naturally classified by its relaxation-rate spectrum than by its excitation spectrum alone.
More broadly, the TDOS framework elevates dynamical spectral organization to a fundamental organizing principle of correlated quantum matter. Memory-dominated criticality represents a universality class distinct from traditional dissipative field theories, with potential implications for non-Fermi liquids, glassy systems, and self-organized critical dynamics.
@@
A dense reservoir of near-marginal slow modes reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$. From this viewpoint, certain high-pressure hydrides or correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat TDOS regime.
The present theory therefore suggests a complementary design principle for superconductivity beyond material-specific glue engineering: stabilization of dynamical phases that support extensive slow-mode continua [27]. Control of electronic frustration, dimensionality, feedback mechanisms, or proximity to dynamical criticality may be as crucial as increasing microscopic attraction itself. Room-temperature superconductivity, in this perspective, corresponds to a dynamical phase characterized by maximal infrared memory rather than simply by extreme coupling strength.
@@
The author acknowledges the support of ChatGPT (GPT-5, OpenAI) for assistance in literature review and conceptual structuring during early development.
@@ Appendix A Appendix A: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
In this appendix we derive how a continuum of relaxation (Ornstein–Uhlenbeck) modes generates a nonlocal-in-time memory kernel and an associated self-energy $\Sigma(\omega)$ upon Gaussian functional integration in the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) formalism. The derivation is standard in structure (Gaussian path integration), but we present it in a form tailored to our “slow-mode reservoir” interpretation.
@@
(VI) Frequency-domain self-energy $\Sigma(\omega)$ from the memory kernel: To connect (S52) to the standard response-function form, we Fourier transform using
@@
(Our convention is $\Sigma_{R}(\omega)\equiv K(\omega)$ so that $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma_{R}(\omega)$; other sign conventions for the self-energy appear in the literature.)
@@
(VII) Connection to TDOS (relaxation-rate density of states): If the couplings are sufficiently structureless over the low-$\lambda$ sector, one may absorb $g(\lambda)^{2}$ into an effective spectral density. Defining
$\mathcal{J}(\lambda)\equiv g(\lambda)^{2},$ (S61)
we can write
$\Sigma_{R}(\omega)=\int d\lambda\;\frac{\mathcal{J}(\lambda)}{\lambda-i\omega}.$ (S62)
Equivalently, starting from a discrete set of modes labeled by $\alpha$,
@@
the continuum limit gives
$\Sigma_{R}(\omega)$ $=\int d\lambda\;\rho(\lambda)\,\frac{\overline{g^{2}}(\lambda)}{\lambda-i\omega}\equiv\int d\lambda\;\frac{\mathcal{J}(\lambda)}{\lambda-i\omega},$
$\qquad\mathcal{J}(\lambda)\equiv\rho(\lambda)\,\overline{g^{2}}(\lambda),$ (S64)
where $\rho(\lambda)$ is the relaxation-rate density of states and $\overline{g^{2}}(\lambda)$ denotes the mean squared coupling of modes with rate $\lambda$.
In our framework the natural object is the time-scale density of states $\rho(\lambda)$ of relaxation rates. In the simplest case $\overline{g^{2}}(\lambda)$ varies weakly over the slow sector so that $\mathcal{J}(\lambda)\propto\rho(\lambda)$ in the infrared, and therefore
$\Sigma_{R}(\omega)\propto\int d\lambda\;\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S65)
This makes explicit, at the level of a controlled Gaussian MSRJD integration, why the relaxation-rate spectrum provides the natural basis for the effective dynamics and response of the collective field.
(VIII) Coarse-grained field extension: We promote the collective coordinate to a coarse-grained field $\phi(\mathbf{r},t)$. At the Gaussian level the MSRJD action reads
@@
$\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega)$ (S68)
(up to the conventional factor $\Gamma$ and the $i\omega$ term in the retarded response).
(IX) Summary of the exact Gaussian result: Starting from the coupled Langevin system (S17)–(S18), the MSRJD action (S24)–(S25) is Gaussian in the reservoir variables. Performing the functional integration over $\{X_{\lambda},\tilde{X}_{\lambda}\}$ yields the exact nonlocal effective action (S52) with a causal memory kernel (S53). In frequency space the kernel takes the spectral form (S58), which is naturally identified with the retarded self-energy (S60). This derivation is purely algebraic and involves no approximation beyond standard boundary-condition assumptions required for integration by parts.
@@ Appendix C Appendix C: Exact emergence of a memory kernel from integrating out an OU reservoir
With this definition, Eq. (LABEL:eq:phi_substituted_app) can be written as
@@
(IV) TDOS form and Laplace-transform structure: The memory kernel identified above depends only on the combination $g(\lambda)^{2}$ that weights the contribution of each relaxational mode to the collective dynamics. It is therefore natural to introduce the TDOS as the effective coupling-weighted spectral density of relaxation rates:
$\rho(\lambda)\equiv g(\lambda)^{2}.$ (S86)
More generally, in a microscopic description with discrete relaxational modes of density $\nu(\lambda)$ and coupling amplitudes $g_{\lambda}$, the kernel involves the combination $\nu(\lambda)|g_{\lambda}|^{2}$. For notational simplicity, we absorb any slowly varying factors into the definition of $\rho(\lambda)$. Thus $\rho(\lambda)$ should be understood as the effective spectral weight governing how strongly relaxation modes at scale $\lambda$ influence the collective coordinate $\phi$.
With this definition, the memory kernel becomes
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},\qquad t\geq 0.$ (S87)
Equation (S87) shows that the memory kernel is precisely the Laplace transform of the TDOS (with ultraviolet cutoff $\Lambda$). In this sense the TDOS is the fundamental spectral object of the theory, while the time-domain memory kernel is its direct transform. The infrared dynamical behavior of the collective coordinate is therefore completely determined by the low-$\lambda$ structure of $\rho(\lambda)$.
@@
Appendix D Appendix D: Renormalization-group flow of the Cooper channel in conventional and memory-dominated regimes
In this Appendix we derive explicitly the infrared renormalization-group (RG) flow equations for the effective pairing interaction in the Cooper channel. We contrast the conventional Eliashberg regime with the memory-dominated regime generated by a flat relaxation-rate density of states. The derivation follows from shell integration of the linearized gap equation and clarifies the origin of marginal versus relevant superconducting instabilities.
Fermionic Matsubara frequencies are denoted by $\omega$ and $\omega^{\prime}$, and $\Omega\equiv\omega-\omega^{\prime}$ is the bosonic transfer frequency entering the pairing kernel $g(\Omega)$.
(I) Linearized gap equation and RG shell construction: At low energies the linearized gap equation takes the schematic form
$\Delta(\omega)=\int_{T}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\,g(\omega-\omega^{\prime})\,\Delta(\omega^{\prime}),$ (S103)
where $\Lambda$ is the ultraviolet cutoff and $T$ provides the infrared cutoff. The factor $1/|\omega^{\prime}|$ originates from the fermionic Cooper bubble. The effective pairing kernel $g(\Omega)$ is generated by collective fluctuations.
The RG procedure consists of integrating out a thin frequency shell $\omega^{\prime}\in[\Lambda/b,\Lambda]$ with $b>1$ and absorbing its contribution into a renormalized coupling $g\rightarrow g+dg$. Defining the RG scale $l=\ln b$, the flow equation follows from the shell contribution.
(II) Conventional Eliashberg regime: Near metallic quantum criticality with Ohmic Landau damping, the bosonic propagator obeys
$D(\Omega)\sim\frac{1}{|\Omega|},$ (S104)
so that the pairing kernel scales as
$g(\Omega)\sim g\,D(\Omega)\sim\frac{g}{|\Omega|},$ (S105)
where $g$ denotes the dimensionless pairing interaction.
The leading shell correction arises from the ladder diagram and is proportional to
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}.$ (S106)
The shell integral is elementary:
$\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{\omega^{\prime}}=\ln\Lambda-\ln(\Lambda/b)=\ln b\equiv dl.$ (S107)
Hence the RG equation takes the marginal form
$\frac{dg}{dl}=C\,g^{2},$ (S108)
with $C$ a positive numerical constant.
The solution is
$g(l)=\frac{g_{0}}{1-Cg_{0}l}\simeq g_{0}+Cg_{0}^{2}l,$ (S109)
demonstrating logarithmic enhancement characteristic of a marginal superconducting instability.
(III) Memory-dominated regime: In the memory-dominated regime derived in Sec. II, the collective susceptibility behaves as
$\chi(\Omega)\sim\frac{1}{\rho_{0}\ln(\Lambda/|\Omega|)},$ (S110)
reflecting the flat TDOS of slow relaxation modes.
The pairing kernel therefore scales as
$g(\Omega)\sim\frac{g}{\ln(\Lambda/|\Omega|)}.$ (S111)
The shell correction becomes
$dg\propto g^{2}\int_{\Lambda/b}^{\Lambda}\frac{d\omega^{\prime}}{|\omega^{\prime}|}\frac{1}{\ln(\Lambda/\omega^{\prime})}.$ (S112)
Introducing the variable
$u=\ln\frac{\Lambda}{\omega^{\prime}},\qquad du=-\frac{d\omega^{\prime}}{\omega^{\prime}},$ (S113)
the integration limits transform as
$\omega^{\prime}=\Lambda\Rightarrow u=0,\qquad\omega^{\prime}=\Lambda/b\Rightarrow u=\ln b\equiv dl.$ (S114)
Equation (S112) becomes
$dg\propto g^{2}\int_{0}^{dl}\frac{du}{u}.$ (S115)
The apparent divergence at $u\to 0$ is cut off at a microscopic scale $u_{\min}=O(1)$ marking the onset of the logarithmic TDOS regime. The shell contribution reduces to
$dg=\alpha\,g\,dl,$ (S116)
where $\alpha>0$ depends on the spectral weight $\rho_{0}$ and the onset scale of the memory-dominated regime.
The RG equation is therefore
$\frac{dg}{dl}=\alpha\,g,$ (S117)
with solution
$g(l)=g_{0}e^{\alpha l}=g_{0}b^{\alpha}.$ (S118)
The pairing interaction thus grows as a power law under infrared coarse graining, signaling a relevant superconducting instability.
(IV) Physical origin of marginal versus relevant behavior: The qualitative difference between Eqs. (S108) and (S117) originates from the infrared structure of the pairing kernel.
In the Eliashberg regime, each logarithmic shell contributes independently, leading to cumulative logarithmic growth. In contrast, the memory-dominated kernel generated by a flat TDOS correlates fluctuations across all time scales, producing linear amplification of the effective interaction under RG.
Consequently, even moderate intrinsic electronic pairing scales are dynamically amplified into a runaway superconducting instability once a slow-mode reservoir is formed.
Summary. Ohmic Landau damping yields a marginal Cooper-channel flow $dg/dl\propto g^{2}$, producing only logarithmic enhancement. A flat relaxation-rate TDOS instead generates a memory-dominated kernel that renders the pairing interaction relevant, with $dg/dl\propto g$ and power-law growth. This dynamical reorganization of infrared spectral weight underlies the algebraic superconducting instability central to the present theory.
Appendix E Appendix E: Finite-temperature regulation of the infrared divergence
@@
Appendix F Appendix F: Spectral function generated by memory–dominated self–energy
@@
(Throughout we use $\omega$ for real frequency; Matsubara frequencies are denoted separately in Appendix D.)
@@
(VII) Physical interpretation: Unlike conventional overdamped criticality, where $\mathrm{Im}\,\Sigma^{R}(\omega)$ vanishes linearly with $|\omega|$, the memory–dominated regime exhibits frequency–independent damping ($\propto\mathrm{sgn}(\omega)$) accompanied by a logarithmically singular reactive part. The resulting spectral function is therefore a broad continuum with slowly varying infrared tails rather than a sharp Lorentzian quasiparticle peak.
Equation (S139) (or equivalently the collapse form (S145)) provides a directly testable dynamical signature of the slow–mode reservoir and the associated TDOS universality class.

### v3

@@ Abstract
From this perspective Hamiltonian energy eigenmodes describe conservative reversible dynamics, whereas relaxation spectra encode the irreversible organization governing collective behavior near criticality. Strongly correlated quantum matter near dynamical critical points may thus be more naturally classified by its relaxation-rate spectrum than by its excitation spectrum alone.
Temporal scaling and dynamical criticality.
An important conceptual aspect of the present framework concerns the role of temporal scaling in dynamical criticality. In conventional renormalization-group treatments of equilibrium critical phenomena, spatial scale invariance provides the primary structure, while temporal scaling emerges indirectly through the dynamic exponent $z$, relating relaxation time and correlation length via $\tau\sim\xi^{z}$. The Hertz–Millis theory of quantum critical metals follows a similar philosophy, in which spatial critical fluctuations generate an effective dynamical scaling once fermionic degrees of freedom are integrated out.
In contrast, the relaxation-spectrum formulation adopted here emphasizes the infrared organization of time scales themselves. When the density of relaxation rates develops a finite limit $\rho(\lambda\to 0)=\rho_{0}$, the system possesses a continuum of slow modes extending to arbitrarily long time scales. The resulting dynamics naturally exhibits scale-free temporal behavior, including long-time memory kernels $K(t)\sim 1/t$ and logarithmic infrared response functions. Scale-free temporal dynamics therefore arises directly from the structure of the relaxation spectrum rather than secondarily from spatial criticality. This perspective is closely related to the scale-free lifetime statistics observed in self-organized critical systems and suggests that infrared relaxation spectra provide a natural language for describing dynamical critical phenomena.
@@
A dense reservoir of near-marginal slow modes reorganizes the Cooper channel, converting moderate intrinsic electronic pairing into an algebraically amplified instability with $T_{c}\sim\rho_{0}E_{\rm pair}$. From this viewpoint certain high-pressure hydrides or correlated materials may fail to reach extreme transition temperatures not because pairing is weak, but because their collective spectra do not self-organize into a flat-TDOS regime.
The present theory therefore suggests a complementary design principle for superconductivity beyond material-specific glue engineering: stabilization of dynamical phases that support extensive slow-mode continua [27]. Control of electronic frustration, dimensionality, feedback mechanisms, or proximity to dynamical criticality may be as crucial as increasing microscopic attraction itself. In this perspective high-temperature superconductivity corresponds to a dynamical phase characterized by maximal infrared memory rather than simply by extreme coupling strength.
@@
The author acknowledges the support of ChatGPT (GPT-5, OpenAI) for assistance in literature review and conceptual structuring during development.
@@ Appendix A Appendix A: Relaxation-rate spectra as the fundamental dynamical basis of collective dynamics
In this appendix we show how a continuum of relaxational (Ornstein–Uhlenbeck) modes generates a nonlocal memory kernel and the associated frequency-dependent self-energy $\Sigma(\omega)$ through Gaussian functional integration within the Martin–Siggia–Rose–Janssen–De Dominicis (MSRJD) formalism. The derivation follows the standard structure of Gaussian path integration but is formulated in a way adapted to the slow-mode reservoir perspective developed in this work.
The construction should be understood as an effective infrared description of collective dynamics rather than a microscopic derivation from an explicit fermionic Hamiltonian. At the microscopic level one may introduce collective Hubbard–Stratonovich fields and integrate out fast fermionic degrees of freedom, in the spirit of Hertz-type constructions, which generally produces a nonlocal dynamical kernel for the collective sector. Instead of evaluating this kernel microscopically, we represent its infrared structure through a continuum of relaxational modes.
Within this coarse-grained description the relaxational reservoir is not an additional assumption but a spectral representation of a causal memory kernel. Integrating out the reservoir therefore yields a retarded dynamical correction and a frequency-dependent self-energy for the collective field.
Such a formulation is particularly natural in strongly correlated systems where the infrared dynamics is governed not by a single soft mode but by a broad spectrum of slow relaxation processes. In this regime the collective response is controlled by the spectral weight distribution of these modes, leading naturally to the TDOS-based description employed throughout this work.
@@
(VII) Frequency-domain self-energy $\Sigma(\omega)$ from the memory kernel: To connect (S52) to the standard response-function form, we Fourier transform using
@@
Our convention is $\Sigma_{R}(\omega)\equiv K(\omega)$ so that $G_{R}^{-1}(\omega)=-i\omega+r+\Sigma_{R}(\omega)$; other sign conventions for the self-energy appear in the literature.
@@
(VIII) Connection to TDOS (relaxation-rate spectral weight): The spectral representation derived above shows that each relaxational mode contributes to the retarded self-energy with pole $(\lambda_{\alpha}-i\omega)^{-1}$ weighted by the square of its overlap with the collective field. Starting from the discrete mode expansion,
@@
it is therefore natural to define the TDOS directly as the coupling-weighted spectral distribution of relaxation rates,
$\rho(\lambda)\equiv\sum_{\alpha}g_{\alpha}^{2}\,\delta(\lambda-\lambda_{\alpha}).$ (S62)
With this definition, the continuum form of the self-energy becomes
$\Sigma_{R}(\omega)=\int d\lambda\;\frac{\rho(\lambda)}{\lambda-i\omega}.$ (S63)
In this sense, $\rho(\lambda)$ plays a role directly analogous to the Lehmann spectral function in Hamiltonian many-body theory. For an operator $O$ with energy eigenstates $|n\rangle$, the usual spectral representation involves weights $|\langle 0|O|n\rangle|^{2}$ multiplying poles at energies $E_{n}$. Here, by contrast, the relevant poles are located at the relaxation rates $\lambda_{\alpha}$, and the corresponding spectral weights are the squared overlap amplitudes $g_{\alpha}^{2}$ between the collective field and the relaxation eigenmodes.
Thus the TDOS should be understood not merely as a bare counting density of modes, but as the physically observable spectral weight distribution of the relaxation spectrum for the collective field under consideration. Equation (S63) makes explicit, at the level of a controlled Gaussian MSRJD integration, why the relaxation-rate spectrum provides the natural basis for the effective dynamics and response of the collective field.
(IX) Coarse-grained field extension: We promote the collective coordinate to a coarse-grained field $\phi(\mathbf{r},t)$. At the Gaussian level the MSRJD action reads
@@
$\chi^{-1}(\omega,\mathbf{q})=r+cq^{2}+\Sigma(\omega),$ (S66)
up to the conventional factor $\Gamma$ and the $i\omega$ term in the retarded response.
Summary of the exact Gaussian result. Starting from the coupled Langevin system (S17)–(S18), the MSRJD action (S24)–(S25) is Gaussian in the reservoir variables. Performing the functional integration over $\{X_{\lambda},\tilde{X}_{\lambda}\}$ yields the exact nonlocal effective action (S52) with a causal memory kernel (S53). In frequency space the kernel takes the spectral form (S58), which is naturally identified with the retarded self-energy (S60). This derivation is purely algebraic and involves no approximation beyond standard boundary-condition assumptions required for integration by parts.
@@ Appendix C Appendix C: Exact emergence of a memory kernel from integrating out an OU reservoir
With this definition, Eq. (LABEL:eq:kernel_identification_step_app) can be written as
@@
(IV) TDOS form and Laplace-transform structure:
As derived in Appendix B, integrating out the continuum of relaxation modes produces a retarded self-energy of the form (S63), where the time-scale density of states is defined as the spectral weight distribution of relaxation eigenmodes,
$\rho(\lambda)=\sum_{\alpha}g_{\alpha}^{2}\,\delta(\lambda-\lambda_{\alpha}).$ (S84)
Here $g_{\alpha}$ represents the overlap amplitude between the collective field and the relaxation eigenmode $\alpha$, so that $g_{\alpha}^{2}$ plays the role of the spectral weight of that mode. In this sense the TDOS is directly analogous to the Lehmann spectral function of Hamiltonian many-body theory, but defined for the relaxation spectrum of the Liouvillian dynamics.
With this definition the memory kernel introduced above can be written in the compact form
$K(t)=\int_{0}^{\Lambda}d\lambda\;\rho(\lambda)\,e^{-\lambda t},\qquad t\geq 0,$ (S85)
showing that the memory kernel is the Laplace transform of the TDOS. Consequently the long-time dynamics of the collective field is fully controlled by the infrared structure of $\rho(\lambda)$.
@@
Appendix D Appendix D: Finite-temperature regulation of the infrared divergence
@@
Figure S1: Experimental diagnostics distinguishing the memory-dominated TDOS regime from Hertz–Millis (HM) dynamics. (a) Absolute spectral line shape at $q=0$: the TDOS theory yields a broad continuum, whereas HM produces a sharper peak. (b) Momentum dependence of the TDOS spectral function: varying $q$ shifts the hump through $\Delta_{q}=r+cq^{2}$. (c) TDOS prediction (collapse): when plotted versus the scaling variable $u=\Delta_{q}/\rho_{0}+\ln(\Lambda/|\omega|)$, the TDOS spectra collapse onto a universal curve, $A(\omega,q)=\frac{\pi}{\rho_{0}}\,[u^{2}+(\pi/2)^{2}]^{-1}$. (d) Control test: HM spectra do not exhibit the same log-shift collapse. (e) Finite-temperature regulation implemented by $|\omega|\rightarrow\max\{|\omega|,cT\}$ yields infrared saturation. (f) Corresponding $T$-diagnostic at $q=0$: $A_{T}(\omega\!\to\!0)$ varies systematically with $\ln(\Lambda/T)$, providing a practical fitting protocol. All curves are shown in arbitrary units with consistent prefactor conventions.
Appendix E Appendix E: Spectral function generated by memory–dominated self–energy
@@
Throughout we use $\omega$ for real frequency; Matsubara frequencies are denoted separately in Appendix D.
@@
Physical interpretation. Unlike conventional overdamped criticality, where $\mathrm{Im}\,\Sigma^{R}(\omega)$ vanishes linearly with $|\omega|$, the memory–dominated regime exhibits frequency–independent damping ($\propto\mathrm{sgn}(\omega)$) accompanied by a logarithmically singular reactive part. The resulting spectral function is therefore a broad continuum with slowly varying infrared tails rather than a sharp Lorentzian quasiparticle peak.
(VII) Experimental diagnostics and scaling tests: Figure S1 summarizes a set of practical diagnostics that distinguish the memory–dominated TDOS regime from conventional overdamped critical dynamics such as the Hertz–Millis (HM) theory.
Panel (a) compares the absolute spectral line shapes at $q=0$ predicted by the two frameworks. The HM theory produces a relatively narrow Lorentzian-like peak, reflecting dynamics controlled by a small number of overdamped bosonic modes. In contrast, Eq. (S121) yields a much broader continuum with slowly varying logarithmic tails, arising from the collective response of an extensive continuum of slow relaxation modes.
Panel (b) illustrates the predicted momentum dependence of the TDOS spectral function. Because the tuning parameter enters through $\Delta_{\mathbf{q}}=r+cq^{2}$, varying the momentum effectively shifts the spectral hump according to the scale $|\omega_{*}|=\Lambda\exp(\Delta_{\mathbf{q}}/\rho_{0})$ derived in Eq. (S125). This produces a characteristic logarithmic displacement of the hump position as $q$ changes.
Panels (c) and (d) demonstrate the most direct experimental test of the TDOS framework. Introducing the scaling variable
$u(\omega,\mathbf{q})=\frac{\Delta_{\mathbf{q}}}{\rho_{0}}+\ln\!\left(\frac{\Lambda}{|\omega|}\right),$ (S130)
all TDOS spectra collapse onto the universal curve (S127). This collapse reflects the fact that the full line shape depends only on the combined variable $u$. By contrast, spectra generated by the conventional Hertz–Millis form do not collapse under the same transformation, providing a clear diagnostic distinction.
Panel (e) illustrates the finite–temperature regulation discussed in Eq. (S129). Replacing $|\omega|$ by $\max\{|\omega|,cT\}$ cuts off the logarithmic infrared growth and produces saturation of the low–frequency spectral weight. The resulting spectra evolve smoothly from the zero–temperature form to a temperature–dominated plateau.
Finally, panel (f) shows a convenient experimental diagnostic at $q\simeq 0$. In the regime $|\omega|\ll T$, Eq. (S129) predicts that the infrared spectral weight varies systematically with $\ln(\Lambda/T)$. Measuring $A_{T}(\omega\!\rightarrow 0)$ as a function of temperature therefore provides a direct way to extract the TDOS parameter $\rho_{0}$ and test the predicted logarithmic scaling.
Taken together, these diagnostics provide a practical experimental protocol for identifying the memory–dominated TDOS universality class. In particular, the logarithmic peak shift and the universal collapse of line shapes constitute strong signatures of collective dynamics governed by an extensive slow–mode reservoir rather than by a single overdamped critical mode.
Appendix F Appendix F: Possible connections between flat TDOS and scale–free dynamical phenomena
In this Appendix we briefly discuss a broader dynamical perspective suggested by the relaxation–spectrum framework developed in the main text. The central quantity of the theory is the density of relaxation rates associated with the Liouvillian generator of the coarse–grained dynamics.
Near dynamical critical regimes, the spectrum of relaxation rates naturally develops a dense continuum extending toward vanishing decay rates. In particular, an important limiting case is a finite infrared density of slow modes,
$\rho(\lambda\rightarrow 0)=\rho_{0}\neq 0.$ (S131)
This condition corresponds to a situation in which relaxation processes exist across arbitrarily long time scales. As a result, the system cannot be characterized by a single relaxation time, and long–time memory effects emerge naturally.
(I) Long–time memory kernel: The dynamical kernel associated with the relaxation spectrum can be written as
$K(t)=\int_{0}^{\Lambda}d\lambda\,\rho(\lambda)\,e^{-\lambda t},$ (S132)
where $\Lambda$ denotes the ultraviolet cutoff of the relaxation spectrum.
If the TDOS approaches a constant for $\lambda\ll\Lambda$, $\rho(\lambda)\approx\rho_{0}$, the long–time behavior becomes
$K(t)\approx\rho_{0}\int_{0}^{\Lambda}d\lambda\,e^{-\lambda t}=\frac{\rho_{0}}{t}\left(1-e^{-\Lambda t}\right).$ (S133)
For $t\gg\Lambda^{-1}$ one therefore obtains the algebraic decay
$K(t)\sim\frac{\rho_{0}}{t},$ (S134)
indicating the emergence of long–time memory.
(II) Relation to scale–free relaxation statistics: Distributions of relaxation times with power–law statistics are widely encountered in complex systems, including models of self–organized criticality (SOC). In many SOC models the lifetime distribution of events takes the approximate form
$P(\tau)\sim\tau^{-2}.$ (S135)
Using the relation between lifetime and relaxation rate $\lambda=1/\tau$, one finds
$\rho(\lambda)=P(\tau)\left|\frac{d\tau}{d\lambda}\right|=P(1/\lambda)\frac{1}{\lambda^{2}}.$ (S136)
Substituting $P(\tau)\sim\tau^{-2}$ yields
$\rho(\lambda)\sim{\rm const}.$ (S137)
Thus a flat infrared TDOS corresponds to a relaxation spectrum that is consistent with scale–free lifetime statistics. This observation suggests a possible dynamical connection between the TDOS framework and the emergence of SOC–type temporal scaling.
(III) Implications for low–frequency noise: A broad distribution of relaxation times is also known to generate 1/f1/f–type noise spectra. Within the present framework the noise spectrum can be written as
$S(\omega)\sim\int d\lambda\,\frac{\rho(\lambda)\lambda}{\lambda^{2}+\omega^{2}}.$ (S138)
When the infrared TDOS approaches a constant, the dominant contribution arises from $\lambda\lesssim\omega$, which yields approximately
$S(\omega)\sim\frac{1}{|\omega|}.$ (S139)
This provides a natural dynamical interpretation of the ubiquitous $1/f$ noise observed in many correlated systems.
(IV) Relation to the infrared spectral function: As shown in Appendix E, a flat TDOS produces a characteristic logarithmic structure in the retarded susceptibility,
$\chi_{R}^{-1}(\omega,q)=\Delta_{q}+\rho_{0}\ln\!\left(\frac{\Lambda}{|\omega|}\right)+i\Gamma\,{\rm sgn}(\omega).$ (S140)
The resulting spectral function
$A(\omega,q)=\frac{2\Gamma}{\left[\Delta_{q}+\rho_{0}\ln(\Lambda/|\omega|)\right]^{2}+\Gamma^{2}}$ (S141)
describes a broad continuum rather than a sharp quasiparticle peak. Such continua have been reported in several strongly correlated systems, including strange metals, although the precise microscopic origin remains an open question.
(V) Consequences for superconducting pairing: In the main text we have shown that the same infrared spectral organization modifies the Cooper channel. The pairing susceptibility takes the form
$\chi_{\Delta}^{-1}(\omega)=\frac{1}{g_{\rm pair}}-\Pi(\omega),$ (S142)
with
$\Pi(\omega)\sim\int d\lambda\frac{\rho(\lambda)}{\lambda^{2}+\omega^{2}}.$ (S143)
For $\rho(\lambda\rightarrow 0)=\rho_{0}$ one finds
$\Pi(\omega)\sim\frac{1}{|\omega|},$ (S144)
which leads to the algebraic transition scale
$T_{c}\sim\rho_{0}E_{\rm pair}.$ (S145)
The slow–mode reservoir therefore acts as a dynamical amplifier of intrinsic electronic pairing correlations.
Summary. These observations suggest that the infrared organization of the relaxation spectrum may provide a common dynamical language for several scale–free phenomena observed in correlated systems. Within this perspective the central object is not a single quasiparticle mode but the entire spectrum of collective relaxation processes. Further investigation of this spectral viewpoint may help clarify the role of dynamical criticality in a wide range of nonequilibrium and strongly correlated systems.